OpenStudy (anonymous):

verify a3+b3 by an activity

OpenStudy (anonymous):

Do you mean \[a^{3} + b^{3}\]

OpenStudy (anonymous):

what do u mean by activity and a3+b3 =(a+b)(a2+b2-ab)

OpenStudy (anonymous):

We know that: \[(a + b)^3 = a^3 + b^3 + 3ab(a + b)\] So, from this let us try to evaluate the value of a^3 + b^3 i.e, \[a^3 + b^3 = (a + b)^3 - 3ab(a + b)\] Solving it further by taking (a + b) common: \[a^3 + b^3 = (a + b)[(a + b)^2 - 3ab]\] \[a^3 + b^3 = (a + b)[a^2 + b^2 + 2ab - 3ab]\] Therefore, the desired value is: \[a^3 + b^3 = (a + b)[a^2 + b^2 - ab]\] That is how you can prove it...